# AdaGC: Improving Training Stability for Large Language Model Pretraining

### v1

@@ Abstract
Large Language Models (LLMs) face increasing loss spikes during scaling, undermining training stability and final performance. While gradient clipping mitigates this issue, traditional global approaches poorly handle parameter-specific gradient variations and decaying gradient norms. We propose AdaGC, an adaptive gradient clipping framework that automatically adjusts local thresholds per parameter through exponential moving average of gradient norms. Theoretical analysis proves AdaGC’s convergence under non-convex conditions. Extensive experiments demonstrate significant improvements: On Llama-2 7B/13B, AdaGC completely eliminates loss spikes while reducing WikiText perplexity by 3.5% (+0.14pp LAMBADA accuracy) for 7B and achieving 0.65% lower training loss with 1.47% reduced validation perplexity for 13B compared to global clipping. For CLIP ViT-Base, AdaGC converges 25% faster than StableAdamW with full spike elimination. The method shows universal effectiveness across architectures (Llama-2 7B/13B) and modalities (CLIP), with successful integration into diverse optimizers like AdamW and Lion. Source code will be released on GitHub.
@@ 1 Introduction
Large Language Models (LLMs) (Brown et al., 2020; Chowdhery et al., 2023; Touvron et al., 2023b) have become a cornerstone of modern Natural Language Processing (NLP) research, demonstrating outstanding performance in a wide range of language tasks. These models, characterized by their extensive parameter counts and vast training datasets, are designed to capture the subtleties of human language. However, as these models grow in size and complexity, they often encounter a significant training obstacle: the loss spike—a sharp, unexpected increase in training loss that can seriously undermine the training process’s efficiency and effectiveness (Chowdhery et al., 2023; Scao et al., 2022; Takase et al., 2023; Zeng et al., 2022).
Among the various methods for addressing spike issues, gradient clipping is an effective solution. Existing gradient clipping approaches (Pascanu et al., 2013) typically apply a predefined threshold to the gradient norm of the entire neural network. However, we have identified two significant shortcomings with this global threshold method. First, as the gradient magnitudes decrease during training, an excessively small threshold throughout the process can slow down the training speed, while a consistently large threshold fails to effectively mitigate spike issues in the later stages of training. Second, due to the inherent differences between structures in neural networks (Zhang et al., 2024), particularly in transformers, applying a uniform clipping coefficient across all layers does not adequately resolve the problems caused by spikes, as illustrated in the Figure 1. To address these two issues with global gradient clipping, we propose using local gradient clipping along with an adaptive update of the clipping threshold for each parameter.
Specifically, we introduce a novel method called Adaptive Gradient Clipping based on Local Gradient Norm (AdaGC). AdaGC is specifically designed to mitigate loss spike by selectively suppressing abnormal gradients, thereby maintaining training stability and enhancing model performance. At its core, AdaGC employs the Exponential Moving Average (EMA) to accurately track historical gradient norm variations at the tensor level. This enables the precise identification and adaptive clipping of anomalous local gradients, effectively containing the loss spike.
Convergence analysis shows AdaGC maintains the $\mathcal{O}(1/\sqrt{T})$ rate comparable to Adam in non-convex settings, confirming that localized clipping preserves convergence properties. Comprehensive experiments on Llama-2 7B/13B (Touvron et al., 2023b) and CLIP ViT-Base (Radford et al., 2021) demonstrate complete loss spike elimination while improving model convergence (3.5% perplexity reduction, 25% faster convergence than StableAdamW (Wortsman et al., 2023a)), with successful integration into modern optimizers like AdamW (Loshchilov & Hutter, 2017) and Lion (Chen et al., 2024).
In summary, our main contributions are three-fold:
  • Identification of two critical limitations in global gradient clipping methods: inadequate handling of decaying gradient norms and parameter-wise heterogeneity, motivating the need for localized adaptive clipping.
  • Proposal of Adaptive Gradient Clipping (AdaGC), featuring parameter-wise threshold adaptation through exponential moving averages, effectively addressing both temporal gradient decay and spatial parameter variations.
  • Comprehensive empirical validation across architectures (Llama-2 7B/13B) and modalities (CLIP ViT-Base), demonstrating complete loss spike elimination with measurable performance gains, plus convergence analysis confirming comparable rates to standard optimizers.
@@ 2 Related Work
Loss spikes present a significant challenge when scaling models, potentially impeding learning or destabilizing the training process. Various strategies have been proposed to enhance the stability of pre-training in large language models.
Training Strategies: Effective training strategies are pivotal for achieving stability. Adopting the bfloat16 format (Wang & Kanwar, 2019) is a cornerstone for stable training. Using smaller learning rates (Wortsman et al., 2023b; Zhang et al., 2022) also promotes stable training dynamics. To mitigate loss spikes, PaLM (Chowdhery et al., 2023) resumes training from a checkpoint approximately 100 steps before the spike and bypasses 200-500 data batches to avoid those preceding and during the spike. Varying the sequence length (Li et al., 2022) is also effective for stabilizing the pre-training of large language models.
Model Architecture: Advanced architectures are crucial for stability in training large-scale models. Research shows that Pre-LN Transformers offer enhanced stability over Post-LN Transformers (Xiong et al., 2020; Vaswani et al., 2017; Liu et al., 2020), both theoretically and practically. Thus, contemporary studies predominantly employ Pre-LN Transformers for building large-scale models. Techniques such as NormHead (Yang et al., 2023), which normalizes output embeddings, and RMSNorm (Zhang & Sennrich, 2019), as reported in Llama (Touvron et al., 2023a), contribute to stable training. Applying layer normalization to the embedding layer (Scao et al., 2022), normalizing the input vector before the standard softmax function (NormSoftmax) (Jiang et al., 2023), and the shrink embedding gradient technique (Zeng et al., 2022) are also beneficial.
Model Initialization: Proper model initialization significantly enhances training stability. Scaling the embedding layer with an appropriate constant (Takase et al., 2023) reinforces LLM pre-training stability. Initializing Transformers with small parameter values (Nguyen & Salazar, 2019) also improves pre-training stability. Some researchers propose that removing layer normalizations in Transformers is feasible with specific initialization methods (Zhang et al., 2019; Huang et al., 2020).
Auxiliary Loss Function: Incorporating an auxiliary loss function alongside conventional language modeling loss enhances training stability. Notable examples include Max-z loss, which benefits the pre-training process (Chowdhery et al., 2023; Wortsman et al., 2023b; Yang et al., 2023).
@@
Gradient/Update Clipping: Gradient and update clipping achieve stability by limiting the magnitude of parameter updates, preventing excessively large weight updates. Global gradient clipping (Pascanu et al., 2013) is prevalent, with innovative approaches like adaptive gradient clipping (Brock et al., 2021) and Clippy (Tang et al., 2023), which use model weights to adjust the clipping threshold. Alternatives like Adafactor (Shazeer & Stern, 2018), StableAdamW (Wortsman et al., 2023a), and LAMB (You et al., 2019) offer update clipping techniques better suited for stability training of large-scale models. Nonetheless, a significant number of loss spikes still occur during the training of large language models, even with the application of these methodologies.
3 Motivations
(a) Temporal threshold adaptation necessity: Fixed thresholds (0.4) prevent early spikes but fail later (10k steps), requiring dynamic adjustment.
(b) Spikes in gradient norms occur at different training steps across parameters, indicating heterogeneous gradient behaviors.
(c) Per-parameter gradient clipping mitigates loss spikes that global clipping cannot, leading to faster convergence.
Figure 1: Empirical motivation for AdaGC: (a) Temporal threshold decay necessitates adaptive clipping, (b) Parameter-specific gradient spikes demand localized control, (c) Fine-grained clipping outperforms global approaches.
In this section, we will introduce some phenomena of gradient clipping that motivate the design of the algorithm.
Adaptivity: The magnitudes of gradients in neural network training exhibit dynamic variations throughout the learning process. This renders fixed gradient clipping thresholds suboptimal for mitigating loss spikes, as demonstrated in Figure 1(a). While a threshold of 0.4 successfully mitigates loss spike during the initial training phase of the GPT-2 345M model, it becomes ineffective against the severe loss spike occurring around 10,000 iterations, where a substantially reduced threshold is required for stabilization. This phenomenon correlates with the observation that optimal clipping thresholds for convergence speed are strongly influenced by recent gradient norms (see corresponding subfigures). Motivated by these findings, we propose an adaptive clipping mechanism using the exponential moving average of gradient magnitudes as the per-step clipping threshold.
Locality: Figure 1(b) reveals that different parameters exhibit divergent gradient behaviors, with spikes occurring at distinct training phases. Using a global clipping threshold can degrade the convergence speed of certain parameters. Therefore, we propose replacing global gradient norm clipping with per-parameter norm clipping, where clipping is applied independently to each parameter. Furthermore, as illustrated in Figure 1(c), we evaluate the effectiveness of per-parameter clipping thresholds categorized by specific ratios on GPT2-345M. The results indicate that, under the same global gradient norm conditions, per-parameter clipping can address spike phenomena that global clipping cannot resolve, and achieve faster convergence speeds.
These observations collectively motivate our two core design principles: adaptive thresholding to address temporal gradient norm decay, and localized control to handle parameter-wise gradient heterogeneity. The dynamic nature of optimal clipping thresholds (Figure 1(a)) necessitates our exponential moving average mechanism for temporal adaptation, while asynchronous parameter gradient spikes (Figure 1(b)) justify per-parameter clipping granularity. Together, these principles form the foundation of AdaGC - adaptively adjusting localized thresholds that respect both the temporal evolution and spatial distribution of gradients.

### v2

# AdaGC: Enhancing LLM Pretraining Stability via Adaptive Gradient Clipping

@@ Abstract
Loss spikes remain a persistent obstacle in large-scale language model pretraining. While previous research has attempted to identify the root cause of loss spikes by investigating individual factors, we observe that, in practice, such spikes are typically triggered by the confluence of heterogeneous factors. Empirically, loss spikes may arise from a combination of data outliers, hardware or transient computational faults, numerical precision issues, and hyperparameter settings. Regardless of the underlying cause, these spikes manifest as unstable optimizer updates, as abnormal gradients contaminate both first- and second-moment states. In this paper, we propose a principled gradient-centric remedy: AdaGC, an adaptive per-tensor gradient clipping scheme that mitigates such contamination by bounding gradient norms relative to a tensor-wise exponential moving average of their historical clipped values. AdaGC is optimizer-agnostic, introduces negligible memory overhead, and reduces communication costs compared to GlobalGC, particularly in hybrid-parallel distributed training environments. Experiments on Llama-2 7B, Mixtral 8×1B, and ERNIE 10B-A1.4B demonstrate that AdaGC robustly eliminates training instabilities, consistently reducing spike scores to zero for all models and improving downstream accuracy over GlobalGC by 1.32%, 1.27%, and 2.48%, respectively. Furthermore, AdaGC seamlessly integrates with optimizers such as Muon and Lion, consistently yielding higher average accuracy and zero spike scores.
@@ 1 Introduction
(a) Mixtral
(b) ERNIE
(c) ERNIE
(d) ERNIE
Figure 1: Reproduced cases of loss spikes and mitigation via resuming. Loss spikes are triggered by (a) increasing $\beta_{2}$ or (b) reducing $\epsilon$ in AdamW, (c) using lower-precision RMSNorm, even under global gradient clipping, and (d) are resolved by resuming due to stochasticity in FlashAttention backward passes.
Table 1: Common sources of loss spikes in LLM pretraining, with typical triggers and practical mitigations.
Source Typical trigger Common mitigation
Data quality Noisy/ill-formed samples Resume and skip data batches
Hardware / transient faults GPU silent errors Fault tolerance and machine replacement
Numerical precision Activation outlier BF16 to FP32 or blockwise quantization
Optimizer / LN hyperparams AdamW/RMSNorm $\epsilon$; AdamW $\beta_{2}$ Careful hyperparameter selection
The rapid scaling of large language models (LLM) has introduced new challenges in pretraining stability, often manifesting as abrupt loss spikes or transient divergences across a wide range of model architectures and data scales (Chowdhery et al., 2023; Touvron et al., 2023; Liu et al., 2024; Team et al., 2025; Baidu-ERNIE-Team, 2025). Despite extensive empirical studies, the fundamental causes of these instabilities remain elusive. Recent research, alongside our own analyses, indicates that loss spikes can arise from a variety of sources, including: (i) data quality issues (Chowdhery et al., 2023); (ii) hardware or transient computational faults (Su, 2025); (iii) variations in numerical precision (for example, FP32 typically offers greater robustness than BF16, whereas FP8 can sometimes enhance stability by suppressing outlier values via implicit quantization (Han, 2024; Liu et al., 2024)); and (iv) the selection of optimizer and layer normalization hyperparameters, such as the $\epsilon$ parameter in RMSNorm or AdamW, and $\beta_{2}$ in AdamW (Ma et al., 2021; Cattaneo & Shigida, 2025; Bai et al., 2025). For instance, we observe that increasing $\beta_{2}$ or decreasing $\epsilon$ in AdamW can trigger loss spikes, whereas increasing the precision of RMSNorm from BF16 to FP32 significantly improves stability. In summary, Table 1 presents the common sources, typical triggers, and practical mitigations for loss spikes identified from both previous studies and our experimental findings, while Figure 1 demonstrates several representative cases we have reproduced.
Although the upstream causes of instability are diverse and often subtle, these events consistently converge at the optimizer level, manifesting as abnormal gradients. Such outlier gradients are incorporated into the optimizer’s first- and second-moment estimates, thereby corrupting parameter updates and propagating instability through subsequent training. Notably, we find that even resuming interrupted training (while keeping the random seed and data unchanged) can mitigate a loss spike, merely due to the stochastic nature of $dQ$, dKdK, and $dV$ in FlashAttention (Dao, 2023) (see Figure 1(d)). This observation further suggests that, in certain model states, even minute numerical differences can trigger a loss spike, with gradient outliers playing a critical role in both the initiation and propagation of instabilities during optimizer state updates.
While the slight stochasticity introduced by FlashAttention can sometimes circumvent a loss spike, repeatedly interrupting and resuming training imposes substantial computational overhead. Given that these instabilities stem from diverse upstream causes but ultimately converge at the optimizer level, our work does not attempt to identify the precise root causes. Instead, we adopt a gradient-centric perspective: irrespective of the initial trigger, loss spikes consistently arise when outlier gradients contaminate the optimizer states. Therefore, by preventing such gradients from entering the first- and second-moment accumulators, we provide a unified and effective strategy to mitigate training instability.
A standard mitigation strategy is global gradient clipping (GlobalGC), which bounds the global $\ell_{2}$ norm of the aggregated gradient. However, this approach is fundamentally mismatched to modern large-scale pretraining in two key respects: (1) Temporal mismatch: The optimal global clipping threshold typically decreases over the course of training; a fixed threshold risks under-clipping in later phases. (2) Spatial mismatch: Gradient statistics and rare spikes vary asynchronously across different parameter tensors, making a single global threshold insufficient—protecting one tensor may under-serve or over-constrain others.
To address these challenges, we introduce Adaptive Gradient Clipping based on Local Gradient Norm (AdaGC): a simple, per-tensor clipping rule that leverages an EMA of each tensor’s historical gradient norm as a reference. Each tensor’s gradient is clipped relative to its own EMA, preventing transient outliers from contaminating the first- and second-moment accumulators and, ultimately, the parameter updates. A brief warm-up period applies global clipping and initializes the EMA to avoid early overestimation. AdaGC is optimizer-agnostic and can be seamlessly integrated with AdamW, Lion, and Muon. Our main contributions are as follows:
  • A unified, gradient-centric perspective: We clarify how loss spikes universally propagate via abnormal gradients polluting optimizer states, irrespective of their origin, motivating intervention at the gradient level prior to moving-average accumulation.
  • An adaptive, per-tensor clipping rule: By tracking each tensor’s gradient norm statistics with an EMA, AdaGC provides both temporal adaptivity and spatial specificity, suppressing outliers while minimally disturbing typical learning dynamics.
  • System efficiency: We analyze computational and communication overhead, showing that AdaGC reduces communication relative to GlobalGC under hybrid parallel distributed training.
[…]
@@ 2 Related Work
Table 2: Comparison of major gradient/update clipping methods for stability. Here, $\boldsymbol{\theta}_{t}$ denotes the model parameters, $\boldsymbol{g}_{t}$ the gradients, $\Delta_{t}$ the optimizer update, $\eta_{t}$ the learning rate, $\lambda_{abs}$ the absolute threshold, and $\lambda_{rel}$ the relative threshold.
Method Algorithm Gradient Update Granularity Threshold Type
GlobalGC (Pascanu et al., 2013) $\min\{1.0,\lambda_{abs}\frac{1}{\|\boldsymbol{g}_{t}\|}\}$ Global Fixed constant
ClipByValue $clamp(-\lambda_{abs},\lambda_{abs})$ Element Fixed constant
AGC (Brock et al., 2021) $\min\{1.0,\lambda_{rel}\frac{\|\boldsymbol{\theta}_{t}\|}{\|\Delta_{t}\|}\}$ Unit Weight $\ell_{2}$ norm
Clippy (Tang et al., 2023) $\min\{1.0,\min(\frac{\lambda_{rel}\|\boldsymbol{\theta}_{t}\|_{\infty}+\lambda_{abs}}{\eta_{t}*\|\Delta_{t}\|_{\infty}})\}$ Tensor Weight $\ell_{\infty}$ norm
AdaGC (ours) $\min\{1.0,\lambda_{rel}\frac{\gamma_{t-1,i}}{\|\boldsymbol{g}_{t,i}\|}\}$ Tensor EMA of gradient norm
$\gamma_{t,i}=\beta\gamma_{t-1,i}+(1-\beta)\|\boldsymbol{g}_{t,i}\|$
Stability in large-scale pretraining: Dozens of approaches address instability during large-model pretraining, including: architectural advances (Pre-LN (Xiong et al., 2020), RMSNorm (Zhang & Sennrich, 2019)), careful initialization (Nguyen & Salazar, 2019; Takase et al., 2023; Nishida et al., 2024), auxiliary loss terms (Max-z loss (Yang et al., 2023)). Recent work (OLMo et al., 2024) also explores combining multiple stabilization strategies. These measures improve average stability but do not directly prevent abnormal gradients from corrupting optimizer states.
@@
Gradient/Update Clipping: Gradient and update clipping achieve stability by limiting the magnitude of gradients and parameter updates, preventing excessively large weight updates. Global gradient clipping (Pascanu et al., 2013) is prevalent, with innovative approaches like AGC (Brock et al., 2021) and Clippy (Tang et al., 2023), which use model weights to adjust the clipping threshold. The SPAM (Huang et al., 2025) method stabilizes the model training process by introducing a momentum reset mechanism and an element-wise gradient clipping strategy based on second-moment estimation. Alternatives like Adafactor (Shazeer & Stern, 2018), StableAdamW (Wortsman et al., 2023), and LAMB (You et al., 2019) offer update clipping techniques better suited for stability training of large-scale models. Nonetheless, a significant number of loss spikes still occur during the training of large language models, even with the application of these methodologies. Due to our gradient-centric perspective, we focus our discussion on clipping-based methods. These methods fall into two categories: value-based approaches, which truncate individual gradient components exceeding a predefined limit, and norm-based approaches, which rescale the entire gradient vector only when its overall magnitude exceeds a threshold. AdaGC belongs to the norm-based category, leveraging adaptive per-tensor norm thresholds to stabilize training. For a comparative summary, see Table 2.
Figure 2: Visualization of the gradient norm, first-moment norm, second-moment norm, update norm, loss, and global gradient norm for the embedding of Llama-2 1.3B during warmup phase. Each row represents a different clipping method: the first row is GlobalGC, the second is AGC, the third is Clippy, and the fourth is our AdaGC. The black curve in each plot shows the loss trajectory.
3 Motivation: From Root-Cause Diversity to a Unified Gradient-Centric Remedy
Through a series of experiments (see Figure 1 and Figure 2), we observe that loss spikes encountered under diverse settings consistently coincide with abrupt fluctuations in the gradient norm. Comparative analyses further reveal limitations of existing methods such as GlobalGC, AGC, and Clippy: GlobalGC’s static global threshold cannot detect or suppress localized abnormal gradients, allowing outliers to contaminate optimizer states and trigger instability. AGC and Clippy focus on controlling parameter updates, leaving internal moments vulnerable to large gradient outliers.
As discussed in the Introduction (Section  1), loss spikes typically result from a combination of multiple factors. While the specific triggers may vary, these loss spikes share a common manifestation: abnormally large gradients are incorporated into the optimizer’s moment estimates, leading to unstable updates. Based on these analyses, we propose a unified remedy: regardless of the root cause, instability in large-scale training is best addressed via gradient-centric clipping. Specifically, only localized and adaptive clipping, applied before gradients are integrated into the optimizer’s moment estimates, can effectively constrain the influence of outlier gradients. We thus distill two key principles for loss spike mitigation: (1) Locality: clip gradients for each parameter tensor individually, avoiding the insensitivity of a global threshold; (2) Adaptivity: dynamically adjust each tensor’s clipping threshold, e.g., using an EMA of its recent gradient norms.
@@ 4.1 Preliminaries
Notations. Let $x_{t}\in\mathbb{R}^{d}$ denote a parameter vector where $x_{t}^{j}$ represents its $j$-th coordinate for $j\in[d]$. We write $\nabla_{x}f(x)$ for the gradient of any differentiable function $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$, and use $u^{2}$ and $u/v$ to denote element-wise square and division operations for vectors $u,v\in\mathbb{R}^{d}$. The $\ell_{2}$-norm and $\ell_{\infty}$-norm are denoted by $\|\cdot\|$ and $\|\cdot\|_{\infty}$, respectively. For asymptotic comparisons, we write $f=\mathcal{O}(g)$ if $\exists c>0$ such that $f(x)\leq cg(x)$ for all $x$ in the domain.
$\small\theta_{t}=\theta_{t-1}-\eta_{t}\nabla_{\theta}f(\theta_{t-1},X_{t})$ (1)
To prevent unstable updates from gradient explosions, GlobalGC (Pascanu et al., 2013) modifies the update rule as:
$\small\begin{array}[]{c}\theta_{t}=\theta_{t-1}-\eta_{t}h_{t}\nabla_{\theta}f(\theta_{t-1},X_{t})\\[3.6999pt] \text{ where }h_{t}:=\min\left\{\frac{\lambda_{abs}}{\|\nabla_{\theta}f(\theta_{t-1};X_{t})\|},1.0\right\}\end{array}$ (2)
Here $\lambda_{abs}$ is an absolute clipping threshold requiring careful tuning, and $\eta_{t}$ is the learning rate. Our work focuses on norm-based clipping (scaling entire gradients exceeding $\lambda_{abs}$) rather than value-based clipping (element-wise truncation).
4.2 Adaptive Gradient Clipping based on Local Gradient Norm
This section introduces a novel gradient clipping strategy termed AdaGC, which distinguishes itself by not relying on a global gradient norm. Instead, AdaGC focuses on the local gradient norm of each tensor and utilizes a dynamic adaptive mechanism for gradient clipping. The proposed method employs an EMA mechanism to maintain smoothed estimates of historical gradient norms per tensor, thus enhancing the accuracy of anomalous gradient detection and enabling independent clipping adjustments tailored to each tensor’s specific conditions. EMA is widely used in deep learning, and within AdaGC, it facilitates the balancing of historical and current gradient norms. The formulation is as follows:
$\small\begin{array}[]{c}\boldsymbol{g}_{t,i}\leftarrow h_{t,i}\cdot\boldsymbol{g}_{t,i},\text{where }h_{t,i}=\min\left\{\lambda_{rel}\frac{\gamma_{t-1,i}}{\|\boldsymbol{g}_{t,i}\|},1.0\right\},\\[9.24994pt] \gamma_{t,i}=\beta\gamma_{t-1,i}+(1-\beta)\|\boldsymbol{g}_{t,i}\|.\end{array}$ (3)
Here, $\lambda_{rel}$ is a predefined relative clipping threshold, $\boldsymbol{g}_{t,i}$ represents the gradient of the $i$-th tensor at time step $t$, and $h_{t,i}$ is a clipping function activated when $\|\boldsymbol{g}_{t,i}\|>\lambda_{rel}\cdot\gamma_{t-1,i}$, thereby scaling the gradient norm to $\lambda_{rel}\cdot\gamma_{t-1,i}$. Additionally, $\beta$ is the smoothing coefficient for EMA. We consistently incorporate the clipped gradient norm into the historical observations rather than the pre-clipped values.
Despite its simplicity, AdaGC adaptively adjusts based on the magnitude of each tensor’s gradient norm. Whenever the gradient norm at a current timestep exceeds a predefined range of average norms within a historical window, it effectively suppresses these outlier gradients.
However, during the initial stages of model training (e.g., the first 100 steps), the gradient norms are typically large and fluctuate significantly, indicating a substantial decreasing trend. Direct application of AdaGC during this period could lead to two issues: first, erroneously accumulating the early large gradient norms into the historical values, resulting in compounded errors; second, compared to GlobalGC, AdaGC might delay clipping, thus potentially slowing down the loss reduction. To address these issues, we introduce a hyperparameter $T_{start}$ (default set to 100), representing a warm-up period during which traditional GlobalGC is applied.
Additionally, AdaGC is optimizer-agnostic, can be seamlessly integrated with various optimizers, such as AdamW (Loshchilov & Hutter, 2017), Lion (Chen et al., 2024), Muon (Jordan et al., 2024), enhancing its practicality and flexibility. Algorithm 1 in Appendix A demonstrates its implementation with the AdamW optimizer.
Table 3: Zero-shot accuracy of AdaGC on Llama-2 7B under different hyperparameters.
$\lambda_{rel}$ $\beta$ 0.98 0.985 0.99 0.999
1.03 50.06 50.92 50.95 50.96
1.04 48.88 50.59 51.04 50.76
1.05 51.01 49.95 50.57 50.74
Table 4: Two-shot accuracy of AdaGC on Llama-2 7B under different hyperparameters.
$\lambda_{rel}$ $\beta$ 0.98 0.985 0.99 0.999
1.03 52.31 52.68 53.13 53.42
1.04 52.68 53.01 53.47 52.85
1.05 52.68 52.67 51.96 53.51
4.3 Memory, Computation, and Communication
Memory. As a tensor-wise method, AdaGC maintains an EMA of gradient norms for each parameter tensor, requiring storage of a single 32-bit float (4 bytes) per tensor. For ERNIE models, the total additional memory overhead has complexity of $\mathcal{O}((9+3E)\times L+3)$, where $L$ and $E$ denote the number of transformer layers and experts, respectively. Specifically, this includes four tensors from the attention module per layer, $3\times(1+E)$ tensors from the shared and router experts per layer, and two RMSNorm tensors per layer; plus one tensor each for the embedding layer, the final layer normalization, and the language modeling head. In practice, this added memory footprint is negligible compared to the overall memory requirements of large-scale model training.
Computation. The computational cost of computing $\ell_{2}$ norms is the same for both AdaGC and GlobalGC. The difference is that GlobalGC applies a uniform scaling to all gradients, while AdaGC scales each gradient tensor independently.
Communication. In setups involving data parallelism (DP), tensor parallelism (TP), and pipeline parallelism (PP), GlobalGC requires an all-reduce operation across all DP, TP, and PP groups to aggregate the global norm. In contrast, AdaGC only needs an all-reduce within each TP group to compute per-tensor local norms. This design substantially reduces communication overhead, offering increasing benefits as model and cluster sizes grow.
5 Experiments
5.1 Experimental Setup